Find the vertex of the parabola
fx=x^2-6x-2

Answer by nerdybill(7384)

fx=x^2-6x-2
you could "complete the square":
fx=(x^2-6x)-2
fx=(x^2-6x+9)-2-9
fx=(x^2-6x+9)-11
fx=(x-3)^2+(-11)
so, vertex is at (3, -11)